mikewwe13:

Find the slope of the line below. A. -1 B. -3/4 C. 3/4 D. 1 E. 4/3

mikewwe13:

mikewwe13:

@jhonyy9 @Mercury

jhonyy9:

do you remember on this formula for this question posted by you past 2 or 3 days ?

mikewwe13:

no but i think the answer is B or C

mikewwe13:

what do you think jhonyy

jhonyy9:

what is the equation of a line in standard form ?

jhonyy9:

really ?

jhonyy9:

than what is the slope from there ?

mikewwe13:

(x1, y1)

mikewwe13:

I might be wrong

jhonyy9:

sorry but i dont see these points in the posted graph there i see given other zeros of the line

mikewwe13:

okay yes i see

jhonyy9:

use this y = mx +b where m is the slope

mikewwe13:

m = y2-y1/x2-x1 = 5-8/5-1 = -3/4

mikewwe13:

is this accurate